OpenStudy (kellyspeakslouder):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f of x equals four divided by x. and g of x equals four divided by x

OpenStudy (kellyspeakslouder):

@triciaal

OpenStudy (kellyspeakslouder):

@jim_thompson5910 this is a writing portion

satellite73 (satellite73):

\[f(x)=\frac{4}{x}\] right?

OpenStudy (kellyspeakslouder):

yes

satellite73 (satellite73):

and also \(g(x)=\frac{4}{x}\)

OpenStudy (kellyspeakslouder):

yes

satellite73 (satellite73):

so what you need is to compute \[f(g(x))\] which only takes two steps first \[f(g(x))=f(\frac{4}{x})\]

satellite73 (satellite73):

then since \[f(\heartsuit)=\frac{4}{\heartsuit}\] you get \[\large f(\frac{4}{x})=\frac{4}{\frac{4}{x}}\]

satellite73 (satellite73):

a little algebra will turn \[\frac{4}{\frac{4}{x}}\] in to \(x\), which is what you want

OpenStudy (kellyspeakslouder):

what are the hearts for?

satellite73 (satellite73):

place holders so you can understand why \[\large f(\frac{4}{x})=\frac{4}{\frac{4}{x}}\]

OpenStudy (kellyspeakslouder):

ohh okay :) thank you! Do I put everything you told me from your 3rd message to your 5th message?

satellite73 (satellite73):

yes, i think so

satellite73 (satellite73):

the point is that \[\large f(g(x))=f(\frac{4}{x})=\frac{4}{\frac{4}{x}}=4\times \frac{x}{4}=x\] and since \(f(g(x))=x\) that means they are inverses or rather "the function is its own inverse"

OpenStudy (triciaal):